OpenStudy (surana):

Which of the following is the correct matrix representation of the September Inventory of small and large T-shirts and pants?

OpenStudy (surana):

\[\left[\begin{matrix}4 & 6 \\ 3 & 5\end{matrix}\right]\] is the September Inventory answer I chose.

OpenStudy (surana):

The September inventory itself is in order of T-Shirts, 4 small and 6 large, and 3 small and 5 large for the pants.

OpenStudy (surana):

I just want to check.

OpenStudy (surana):

Other possible answers were \[\left[\begin{matrix}2 & 3 \\ 1 & 4\end{matrix}\right]\] \[\left[\begin{matrix}7 & 8 \\ 5 & 6\end{matrix}\right]\] and \[\left[\begin{matrix}5 & 6 \\ 1 & 4\end{matrix}\right]\]

OpenStudy (surana):

Closed due to inactivity and because I'm fairly certain I'm correct.